OpenStudy (anonymous):

let log P/N = 6 and log M/N = 8. what is the relationship between P ad M?

OpenStudy (anonymous):

Let the log be some base x. P/N = x^6 and M/N = x^8 divide to get P/M = x^-2

OpenStudy (dumbcow):

for example if this is log base 10 , then M/P = 10^2 = 100 M is 100 times bigger than P

OpenStudy (anonymous):

ohhhhh ok then

OpenStudy (anonymous):

soooo technically putting in a equation it would be put together like this ... M = 100P @dumbcow

OpenStudy (dumbcow):

yes if the base is 10 M = x^2 P for any base "x"

OpenStudy (anonymous):

ohhh ok then n_n thank you i understant it @dumbcow